A Bibliography of Font Readability

Karl Berry 135 Center Hill Rd. Plymouth, MA 02360

USA

E-mail: karl@cs.umb.edu (Internet)

26 October 2010 Version 1.33

Title word cross-reference

Acuity [O’R88]. Advances [Pie91]. affect [MBHL91]. American [KF67]. Analysis [KF67]. Angle [MR81].

Character [AKG90, MR81]. Clarity [WNS87]. Close [MS84]. Close-Set [MS84]. Comfort [WNS87]. Computational [KF67]. Congress [Pie91]. Content [BLL91]. Contrast [LRL87].

Day [KF67]. Depends [MR81].

English [KF67]. Eye [O’R88].

Fixed [AKG90]. Frequency [BLL91].

impact [Pie91]. International [Pie91]. IS&T [Pie91].

Legibility [Tin63]. Letters [BLL91]. Lexical [O’R88].

Movements [O’R88].

Non [Pie91]. Non-impact [Pie91]. Normal [LRL87].

October [Pie91].

Pitch [AKG90]. Portland [Pie91]. Present [KF67]. Present-Day [KF67]. Print [Tin63]. Printed [WNS87]. Printing [Pie91]. Psychophysics [LRL87].

readability [MBHL91]. Reading [AKG90, BLL91, LRL87, MR81]. Recognition [O’R88]. Role [LRL87].

Saccade [MR81]. Set [MS84]. Seventh [Pie91]. Size [MR81]. sizes [MBHL91]. small [MBHL91]. Spaces [MR81]. Spatial [BLL91]. Spatial-Frequency [BLL91]. Structure [O’R88]. Study [MS84].

Technologies [Pie91]. Text [MS84, WNS87]. Their [BLL91]. Type [MS84]. typeface [MBHL91]. typographic [MBHL91].

Upon [MR81]. USA [Pie91].

V [LRL87]. Variable [AKG90]. variation [MBHL91]. Vision [LRL87]. Visual [MR81, O’R88].

Word [O’R88]. Words [BLL91].

References

Arditi:JOSAA-7-2011

[AKG90] Aries Arditi, Kenneth Knoblauch, and Ilana Grunwald. Reading with fixed and variable character pitch. J. Opt. Soc. Amer., 7(10):2011– 2015, 1990.

Beckmann:SID91-106

[BLL91] Paul J. Beckmann, Gordon E.

Legge, and Andrew Luebker. Reading: Letters, words, and their spatial-frequency content. SID 91 Digest, pages 106–108, 1991.

Kucera:CAE-67

[KF67] Henry Kuĉera and W. Nelson Francis. Computational Analysis of Present-Day American English. Brown University Press, Providence, RI, USA, 1967.

Legge:VR-27-1165

[LRL87] Gordon E. Legge, Gary S. Rubin, and Andrew Luebker. Psychophysics of reading—V. The role of contrast in normal vision. Vision Res., 27(7):1165–1177, 1987.

Morris:NIPT91-344

[MBHL91] Robert A. Morris, Karl Berry, Kathryn A. Hargreaves, and Dimitrios Liarokapis. How typeface variation and typographic variation affect readability at small sizes. In Pietrowski [Pie91], pages 344–352.

Morrison:PP-30-395

[MR81] Robert E. Morrison and Keith Rayner. Saccade size in reading depends upon character spaces and not visual angle. Perception & Psychophysics, 30(4):395–396, 1981.

Moriarty:AP-69-700

[MS84] Sandra Ernst Moriarty and Edward C. Scheiner. A study of closeset text type. J. of Applied Psychology, 69:700–703, 1984.

ORegan:WMH-88-261

[O’R88] J. Kevin O’Regan. Visual acuity, lexical structure and eye movements in word recognition. In Working Models of Human Perception. 1988.

Pietrowski:NIPT91

[Pie91] Ken Pietrowski, editor. IS&T’s Seventh International Congress on Advances in Non-impact Printing Technologies, October 6–11, Portland, OR, USA., volume 2, 1991.

Tinker:LP-63

[Tin63] Miles A. Tinker. Legibility of Print. Iowa State University, Ames, IA, USA, 1963.

Wilkins:ERG-30-1705

[WNS87] A. J. Wilkins and M. I. Nimmo-Smith. The clarity and comfort of printed text. Ergonomics, 30(12): 1705–1720, 1987.